from a balloon which is at height H= 300m above the ground a stone is left to fall freely . How long will it take for the stone to reach the ground 1) balloon is moving up at velocity v= 5m/s 2) balloon is moving down at velocity v= 5m/s 3) balloon is motionless
1
Expert's answer
2018-03-16T08:04:07-0400
1)
h=-vt+(gt^2)/2
300=-5t+(10t^2)/2
t=8.3 s.

2)
h=vt+(gt^2)/2
300=5t+(10t^2)/2
t=7.3 s.
3)
h=(gt^2)/2
t=√(2h/g)=√((2(300))/10)=7.8 s.